OpenStudy (anonymous):

18. Solve: -10(2x + 1) + 3(4 - x) = 71 A. x = 69 B. x = -3 C. x = 2.9 D. x = 3 19. Solve: 4x - 19 = 7x + 2 A. x = 7 B. x = -1.55 C. no solution D. x = -7 20. Solve: 14x + 10 - 8x = x - 20 A. x = 6 B. x = -6 C. no solution D. x = 4

OpenStudy (anonymous):

18. multiply in the -10 and 3 simplify and solve for x -20x-10 + 12 - 3x= 71 -23x +2 = 71 -23x = 69 x = -3 19. simplify and solve for x 3x = -21 x=-7 20.simplify and solve for x 5x = -30 x = -6

OpenStudy (anonymous):

thank you i suck at math

OpenStudy (anonymous):

np glad to help